OpenStudy (anonymous):

please help!!! Find the surface area of a rectangular prism that has the following dimensions. length: 4.5 in. width: 6 in. height: 12 in.

jimthompson5910 (jim_thompson5910):

Surface Area of a Rectangular Prism = 2*(Length*Width + Length*Height + Width*Height) SA = 2*(LW + LH + WH) SA = 2*(4.5*6 + 4.5*12 + 6*12) SA = 2*(27 + 54 + 72) SA = 2*(153) SA = 306 So the surface area of the rectangular prism is 306 square inches.